OpenStudy (anonymous):

write a polynomial function of leaswt degree with integral coefficients that have the given zeros.: -2,5,-3i

OpenStudy (anonymous):

(x+2)(x-5)(((-x)^2)+9)

OpenStudy (mathmath333):

solutions are given as -2,5,-3i,therefore the polynomial equation will be \(\large\tt \color{black}{=(x+2)(x-5)(x-3i)(x+3i)}\)

OpenStudy (mathmath333):

simplify it now

OpenStudy (mathmath333):

its very easy

OpenStudy (mathmath333):

first lets simplify (x+2)(x-5)

OpenStudy (mathmath333):

(x+2)(x-5) =x^-5x+2x-10 =x^2-3x-10

OpenStudy (mathmath333):

now lets simplify (x-3i)(x+3i) u can simplify it as i did above and use formula here for quick solution (a+b)(a-b)=a^2-b^2 hence (x-3i)(x+3i) =x^-(3i)^2 =x^2-9i^2 =x^2+9..............(as i=\[ i=\sqrt{-1}\] (i^2=-1) now the equation will be (x+2)(x-5)(x-3i)(x+3i) =(x^2-3x-10)(x^2+9) =x^4+9x^2-3x^3-27x-10x^2-90 =x^4-3x^3-x^2-27x-90 so this is ur polynomial function f(x)==x^4-3x^3-x^2-27x-90

OpenStudy (mathmath333):

\(\huge\tt \color{black}{f(x)=x^4-3x^3-x^2-27x-90}\)